OpenStudy (anonymous):

Suppose y varies directly with x, and y = 14 when x = 4. What is the value of y when x = 9

OpenStudy (confluxepic):

\[\huge \color{blue}{Welcome~To~OpenStudy!!!}\]

OpenStudy (xapproachesinfinity):

y varies directly with x men y=kx find k if y=14 and x=4 then we 14=4k ===> k=14/4=7/2 if x=9 we have y=7/2 x 9=63/2